OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). y = 4/3x + 8 y = 4/3x + 3 y = -3/4x + 8 y = -3/4x + 3 Thanks!

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Step 1 change 4x - 3y = -12 into the slope-intercept form

OpenStudy (skullpatrol):

$$\Huge y=mx +b$$

OpenStudy (skullpatrol):

Step 2 using the slope m and the given point (-3, 4) write the answer :)

OpenStudy (skullpatrol):

Answer: y = 4/3x + 8

OpenStudy (anonymous):

Thanks :)